Juanita invests $3,000, at 5% interest, compounded annually for 3 years. Manually calculate the compound interest for this investment.
I understand every problem but the ones like these, Please can you help step me through the process to find the answer? I either skipping a step or something... I can't figure it out
Thanks

Answer by rothauserc(4718)

the formula for compound interest is,
A = P*(1 + r/n)^nt where A is amount, P is principal, r is rate, n is number of times the rate is compounded per year, t is number of years
A = 3000*(1 + .05/1)^(1*3) = 3472.875 approx $3472.88
compound interest is 3472.88 - 3000.00 = $472.88